# Tony Lindeberg

### v5

@@ Abstract
journal: Submitted to Journal of Mathematical Imaging and Vision
@@ 1 Introduction
A practical observation, that one may make, when working with deep learning, is that deep networks may tend to have a preference to computing image representations at very fine scale levels. For example, empirical results indicate that deep networks often tend to perform image classification based on very fine-scale image information, corresponding to the local image texture on the surfaces of objects in the world. Indirect support for such a view may also be taken from then now well-established fact that deep networks may be very sensitive to adversarial perturbations, based on adding deliberately designed noise patterns of very low amplitude to the image data (Szegedy et al. 2013, Moosavi-Dezfooli et al. 2017, Athalye et al. 2018, Baker et al. 2018, Hendrycks et ak. 2021). That observation demonstrates that deep networks may be very strongly influenced by fine-scale structures in the input image. Another observation may be taken from working with deep networks based on using Gaussian derivative kernels as the filter weights. If one designs such a network with complementary training of the scale levels for the Gaussian derivatives, then a common result is that the network will prefer to base its decisions based on receptive fields at rather fine scale levels.
@@
Rey-Otero and Delbracio (2016) assumed that the image data can be regarded as bandlimited, and did then use a Fourier-based approach for performing closed-form Gaussian convolution over a reconstructed Fourier basis, which in that way constitutes a way to eliminate the discretization errors, provided that a correct reconstruction an underlying continuous image, notably before the image acquisition step, can be performed. [^1] For a number of clarifications with respect to an evaluation of what Rey-Otero and Delbracio (2016) refer to as “Lindeberg’s smoothing method” in that work, see Appendix A.10. A very closely related approach, for computing Gaussian convolutions based on a reconstruction of an assumed-to-be bandlimited signal, has also been previously outlined by Åström and Heyden (1997).
@@
For performing such analysis, we will consider basic approaches for discretizing the Gaussian kernel in terms of either pure spatial sampling (the sampled Gaussian kernel) or local integration over each pixel support region (the integrated Gaussian kernel) and compare to the results of a genuinely discrete scale-space theory (the discrete analogue of the Gaussian kernel). After analysing and numerically quantifying the properties of these basic types of discretizations, we will then extend the analysis to discretizations of Gaussian derivatives in terms of either sampled Gaussian derivatives, integrated Gaussian derivatives, and compare to the results of a genuinely discrete theory based on convolutions with the discrete analogue of the Gaussian kernel followed by discrete derivative approximations computed by applying small-support central difference operators to the discrete scale-space representation. We will also extend the analysis to the computation of local directional derivatives, as a basis for filter-bank approaches for receptive fields, based on either the scale-space representation generated by convolution with rotationally symmetric Gaussian kernels, or the affine Gaussian scale space.
@@ 3.3 Sampled Gaussian derivative kernels
Figures 8(a)8(d) and Figures 9(a)9(d) show how the $l_{1}$-norms as well as the spatial spread measures vary as function of the scale parameter, with comparisons to the scale dependencies for the corresponding fully continuous measures.
@@ 3.4 Integrated Gaussian derivative kernels
With this definition, it follows immediately that the contributions to the $l_{1}$-norm of the discrete kernel $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\;s)$ will be equal to the contributions to the $L_{1}$-norm of $g_{x^{\alpha}}(n;\;s)$ over those pixels where the continuous kernel has the same sign over the entire pixel support region. For those pixels where the continuous kernel changes its sign within the support region of the pixel, however, the contributions will be different, thus implying that the contributions to the $l_{1}$-norm of the discrete kernel may be lower than the contributions to the $L_{1}$-norm of the corresponding continuous Gaussian derivative kernel.
Similarly to the previously treated case with the integrated Gaussian kernel, the integrated Gaussian derivative kernels will also imply a certain scale offset, as shown in Figures 9(a)9(d) and Figures 10(a)10(d).
@@ 3.6 Numerical correctness of the derivative estimates
With respective to Gaussian derivative responses to monomials of the form
@@
Figures 6(a)6(a) and Figures 7(a)7(b) show the results of computing the responses of the discrete approximations of Gaussian derivative operators to different monomials in this way, up to order 4. Specifically, Figure 6(a) shows the entity $P_{1,1}(s)$, which in the continuous case should be equal to 1. Figure 6(b) shows the entity $P_{2,2}(s)$, which in the continuous case should be equal to 2. Figure 6(c) shows the entity $P_{3,3}(s)$, which in the continuous case should be equal to 6. Figure 6(d) shows the entity $P_{4,4}(s)$, which in the continuous case should be equal to 24. Figure 7(a) and Figure 7(b) show the entities $P_{3,1}(s)$ and $P_{4,2}(s)$, respectively, which in the continuous case should be equal to zero.
@@ 3.8.1 $l_{1}$-norms of discrete approximations of Gaussian derivative approximation kernels
Figures 8(a)8(d) show the $l_{1}$-norms $\|T_{x^{\alpha}}(\cdot;\,s)\|_{1}$ for the different methods for approximating Gaussian derivative kernels with corresponding discrete approximations for differentiation orders up to 4, together with graphs of the $L_{1}$-norms $\|g_{x^{\alpha}}(\cdot;\,s)\|_{1}$ of the corresponding continuous Gaussian derivative kernels.
@@ 3.8.2 Spatial spread measures
Figures 9(a)9(d) show graphs of the standard-deviation-based spatial spread measure $\sqrt{V(|T_{x^{\alpha}}(\cdot;\;s)|)}$ according to (80), for the main classes of discretizations of Gaussian derivative kernels, together with graphs of the corresponding spatial spread measures computed for continuous Gaussian derivative kernels.
@@
Figures 10(a)10(d) show more detailed visualizations of the deviations between these spatial spread measures and their corresponding ideal values for continuous Gaussian derivative kernels in terms of the spatial spread measure offset $O_{\alpha}(s)$ according to (81), for the different orders of spatial differentiation. The jaggedness of these curves for orders of differentiation greater than one, is due to interactions between the lobes in the derivative approximation kernels and the grid. As can be seen from these graphs, the relative properties of the spatial spread measure offsets for the different discrete approximations to the Gaussian derivative operators differ somewhat, depending on the order of spatial differentiation. We can, however, note that the spatial spread measure offset for the integrated Gaussian derivative kernels is mostly somewhat higher than the spatial spread measure offset for the sampled Gaussian derivative kernels, consistent with the previous observation that the spatial box integration used for defining the integrated Gaussian derivative kernel introduces an additional amount of spatial smoothing in the spatial discretization.
@@ 3.8.3 Cascade smoothing errors
Figures 11(a)11(d) show graphs of the cascade smoothing error $E_{\scriptsize\mbox{cascade}}(T_{x^{\alpha}}(\cdot;\;s))$ according to (82) for the main classes of methods for discretizing Gaussian derivative operators.
@@ 4.4 Measures of scale selection performance
Figure 14: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (107), when applying scale selection from local extrema over scale of the scale-normalized Laplacian response according to (4.3.1) to a set of Gaussian blobs of different size $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (64), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (53), or integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (54). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[0.1,4]$.)
@@
Figure 16: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (107), when applying scale selection from local extrema over scale of the scale-normalized determinant of the Hessian response according to (4.3.1) to a set of Gaussian blobs of different size $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (64), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (53), or integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (54). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[0.1,4]$.)
@@
Selected scales for gradient magnitude scale selection
@@
Figure 18: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (107), when applying scale selection from local extrema over scale of the scale-normalized gradient magnitude response according to (100) to a set of diffuse step edges of different width $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (64), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (53), or integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (54). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[0.1,4]$.)
@@
Selected scales for principal curvature scale selection
@@
Figure 20: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (107), when applying scale selection from local extrema over scale of the scale-normalized principal curvature response according to (105) to a set of diffuse ridges of different width $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (64), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (53), or integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (54). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[0.1,4]$.)
@@ 5.1 Small-support directional derivative approximation masks
If we, on the other hand, base the analysis on the discrete scale-space concept, by which derivative approximations can be computed from the raw discrete scale-space representation, by applying small-support central difference masks, then the amount of computational work can be decreased substantially, since then for each new combination of the orders $m_{1}$ and $m_{2}$ of differentiation, we only need to apply a new small-support discrete filter mask. In the case, when the underlying scale-space representation is based on convolutions the rotationally symmetric Gaussian kernel, we can use the same underlying, once and for all spatially smoothed image as the input for computing filter bank responses for all the possible orientations. In the case, when the underlying scale-space representation is instead based on convolutions with affine Gaussian kernels, we do, of course, have to redo the underlying spatial smoothing operation for each combination of the parameters $\sigma_{1}$, $\sigma_{2}$ and $\varphi$. We can, however, nevertheless reuse the same underlying spatially smoothed image for all the combinations of the orders $m_{1}$ and $m_{2}$ of spatial differentiation.
@@ 5.2 Method for defining discrete directional derivative approximation masks
To define a discrete derivative approximation mask $\delta_{\varphi^{m_{1}}\bot\varphi^{m_{2}}}$, for computing an approximation the directional derivative $L_{\varphi^{m_{1}}\bot\varphi^{m_{2}}}$ from an already smoothed scale-space representation $L$ according to
@@ 6 Summary and conclusions
Regarding the purely spatial smoothing operation, the discrete analogue of the Gaussian kernel stands out as having the best theoretical properties over the entire scale range, from scale levels approaching zero to large scales. The results obtained from the sampled Gaussian kernel may deviate substantially from their continuous counterparts, when the scale parameter $\sigma$ is less than about 0.5 or 0.75. For $\sigma$ greater than about 1, the sampled Gaussian kernel does, on the other hand, lead to numerically very good approximations of results obtained from the corresponding continuous theory.
@@
The fact that the sampled Gaussian derivative kernels for sufficiently large scales lead to the closest approximations of the corresponding fully continuous theory should, however, not preclude from basing the analysis on the discrete analogues of Gaussian derivatives at coarser scales. If necessary, deviations between the results obtained from the discrete analogues of Gaussian derivatives and the corresponding fully continuous theory can, in principle, be compensated for by complementary calibration procedures, or by deriving corresponding genuinely discrete analogues of the relevant entities in the analysis. Additionally, in situations when a larger number of Gaussian derivative responses are to be computed simultaneously, this can be accomplished with substantially higher computational efficiency, if basing the scale-space analysis on the discrete analogue of the Gaussian kernel, which only involves a single spatial smoothing stage of large spatial support, from which each derivative approximation can then be computed using a small-support central difference operator.
As a complement to the presented methodologies of discretizing Gaussian smoothing and Gaussian derivative computations, we have also in Section 5 presented a computationally very efficient ways of computing directional derivatives of different orders and of different orientations, which is highly useful for computing filter bank type responses for different purposes in computer vision. When using the discrete analogue of the Gaussian kernel for smoothing, the presented discrete directional derivative approximation masks can be applied at any scale. If using either sampled Gaussian kernels or integrated Gaussian kernels for spatial smoothing, including extensions of from rotationally symmetric kernels to anisotropic affine Gaussian kernels, the discrete derivative approximation masks can be used, provided that the scale parameter is sufficiently large in relation to the desired accuracy of the resulting numerical approximation.

### v7

@@ Abstract
journal: arXiv preprint
@@ 1 Introduction
A practical observation, that one may make, when working with deep learning, is that deep networks may tend to have a preference to computing image representations at very fine scale levels. For example, empirical results indicate that deep networks often tend to perform image classification based on very fine-scale image information, corresponding to the local image texture on the surfaces of objects in the world. Indirect support for such a view may also be taken from the now well-established fact that deep networks may be very sensitive to adversarial perturbations, based on adding deliberately designed noise patterns of very low amplitude to the image data (Szegedy et al. 2013, Moosavi-Dezfooli et al. 2017, Athalye et al. 2018, Baker et al. 2018, Geirhos et al. 2018, Hermann et ak. 2020, Hendrycks et ak. 2021, Veerabadran et al. 2023). That observation demonstrates that deep networks may be very strongly influenced by fine-scale structures in the input image. Another observation may be taken from working with deep networks based on using Gaussian derivative kernels as the filter weights. If one designs such a network with complementary training of the scale levels for the Gaussian derivatives, then a common result is that the network will prefer to base its decisions based on receptive fields at rather fine scale levels.
@@
Rey-Otero and Delbracio (2016) assumed that the image data can be regarded as bandlimited, and did then use a Fourier-based approach for performing closed-form Gaussian convolution over a reconstructed Fourier basis, which in that way constitutes a way to eliminate the discretization errors, provided that a correct reconstruction of an underlying continuous image, notably before the image acquisition step, can be performed. [^1] For a number of clarifications with respect to an evaluation of what Rey-Otero and Delbracio (2016) refer to as “Lindeberg’s smoothing method” in that work, see Appendix A.10. A very closely related approach, for computing Gaussian convolutions based on a reconstruction of an assumed-to-be bandlimited signal, has also been previously outlined by Åström and Heyden (1997).
@@
For performing such analysis, we will consider basic approaches for discretizing the Gaussian kernel in terms of either pure spatial sampling (the sampled Gaussian kernel) or local integration over each pixel support region (the integrated Gaussian kernel) and compare to the results of a genuinely discrete scale-space theory (the discrete analogue of the Gaussian kernel), see Figure 1 for graphs of such kernels.d After analysing and numerically quantifying the properties of these basic types of discretizations, we will then extend the analysis to discretizations of Gaussian derivatives in terms of either sampled Gaussian derivatives, integrated Gaussian derivatives, and compare to the results of a genuinely discrete theory based on convolutions with the discrete analogue of the Gaussian kernel followed by discrete derivative approximations computed by applying small-support central difference operators to the discrete scale-space representation. We will also extend the analysis to the computation of local directional derivatives, as a basis for filter-bank approaches for receptive fields, based on either the scale-space representation generated by convolution with rotationally symmetric Gaussian kernels, or the affine Gaussian scale space.
@@ 3.3 Sampled Gaussian derivative kernels
Figures 9(a)9(d) and Figures 10(a)10(d) show how the $l_{1}$-norms as well as the spatial spread measures vary as function of the scale parameter, with comparisons to the scale dependencies for the corresponding fully continuous measures.
@@ 3.4 Integrated Gaussian derivative kernels
With this definition, it follows immediately that the contributions to the $l_{1}$-norm of the discrete kernel $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\;s)$ will be equal to the contributions to the $L_{1}$-norm of $g_{x^{\alpha}}(n;\;s)$ over those pixels where the continuous kernel has the same sign over the entire pixel support region. For those pixels where the continuous kernel changes its sign within the support region of the pixel, however, the contributions will be different, thus implying that the contributions to the $l_{1}$-norm of the discrete kernel may be lower than the contributions to the $L_{1}$-norm of the corresponding continuous Gaussian derivative kernel (see Figure 1 for an illustration of such graphs of integrated Gaussian derivative kernels).
Similarly to the previously treated case with the integrated Gaussian kernel, the integrated Gaussian derivative kernels will also imply a certain scale offset, as shown in Figures 10(a)10(d) and Figures 11(a)11(d).
@@ 3.6 Numerical correctness of the derivative estimates
With respect to Gaussian derivative responses to monomials of the form
@@
Figures 7(a)7(d) and Figures 8(a)8(b) show the results of computing the responses of the discrete approximations of Gaussian derivative operators to different monomials in this way, up to order 4. Specifically, Figure 7(a) shows the entity $P_{1,1}(s)$, which in the continuous case should be equal to 1. Figure 7(b) shows the entity $P_{2,2}(s)$, which in the continuous case should be equal to 2. Figure 7(c) shows the entity $P_{3,3}(s)$, which in the continuous case should be equal to 6. Figure 7(d) shows the entity $P_{4,4}(s)$, which in the continuous case should be equal to 24. Figure 8(a) and Figure 8(b) show the entities $P_{3,1}(s)$ and $P_{4,2}(s)$, respectively, which in the continuous case should be equal to zero.
@@ 3.8.1 $l_{1}$-norms of discrete approximations of Gaussian derivative approximation kernels
Figures 9(a)9(d) show the $l_{1}$-norms $\|T_{x^{\alpha}}(\cdot;\,s)\|_{1}$ for the different methods for approximating Gaussian derivative kernels with corresponding discrete approximations for differentiation orders up to 4, together with graphs of the $L_{1}$-norms $\|g_{x^{\alpha}}(\cdot;\,s)\|_{1}$ of the corresponding continuous Gaussian derivative kernels.
@@ 3.8.2 Spatial spread measures
Figures 10(a)10(d) show graphs of the standard-deviation-based spatial spread measure $\sqrt{V(|T_{x^{\alpha}}(\cdot;\;s)|)}$ according to (80), for the main classes of discretizations of Gaussian derivative kernels, together with graphs of the corresponding spatial spread measures computed for continuous Gaussian derivative kernels.
@@
Figures 11(a)11(d) show more detailed visualizations of the deviations between these spatial spread measures and their corresponding ideal values for continuous Gaussian derivative kernels in terms of the spatial spread measure offset $O_{\alpha}(s)$ according to (81), for the different orders of spatial differentiation. The jaggedness of these curves for orders of differentiation greater than one, is due to interactions between the lobes in the derivative approximation kernels and the grid. As can be seen from these graphs, the relative properties of the spatial spread measure offsets for the different discrete approximations to the Gaussian derivative operators differ somewhat, depending on the order of spatial differentiation. We can, however, note that the spatial spread measure offset for the integrated Gaussian derivative kernels is mostly somewhat higher than the spatial spread measure offset for the sampled Gaussian derivative kernels, consistent with the previous observation that the spatial box integration used for defining the integrated Gaussian derivative kernel introduces an additional amount of spatial smoothing in the spatial discretization.
@@ 3.8.3 Cascade smoothing errors
Figures 12(a)12(d) show graphs of the cascade smoothing error $E_{\scriptsize\mbox{cascade}}(T_{x^{\alpha}}(\cdot;\;s))$ according to (82) for the main classes of methods for discretizing Gaussian derivative operators.
@@ 4.4 Measures of scale selection performance
Figure 14: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (107), when applying scale selection from local extrema over scale of the scale-normalized Laplacian response according to (4.3.1) to a set of Gaussian blobs of different size $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (64), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (53), or integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (54). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[1/3,4]$.)
Selected scales for detHessian scale selection
@@
Figure 16: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (107), when applying scale selection from local extrema over scale of the scale-normalized determinant of the Hessian response according to (4.3.1) to a set of Gaussian blobs of different size $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (64), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (53), or integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (54). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[1/3,4]$.)
Selected scales for gradient magnitude scale selection
@@
Figure 18: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (107), when applying scale selection from local extrema over scale of the scale-normalized gradient magnitude response according to (100) to a set of diffuse step edges of different width $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (64), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (53), or integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (54). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[1/3,4]$.)
Selected scales for principal curvature scale selection
@@
Figure 20: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (107), when applying scale selection from local extrema over scale of the scale-normalized principal curvature response according to (105) to a set of diffuse ridges of different width $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (64), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (53), or integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (54). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[1/3,4]$.)
@@ 5.1 Small-support directional derivative approximation masks
If we, on the other hand, base the analysis on the discrete scale-space concept, by which derivative approximations can be computed from the raw discrete scale-space representation, by applying small-support central difference masks, then the amount of computational work can be decreased substantially, since then for each new combination of the orders $m_{1}$ and $m_{2}$ of differentiation, we only need to apply a new small-support discrete filter mask. In the case, when the underlying scale-space representation is based on convolutions with the rotationally symmetric Gaussian kernel, we can use the same underlying, once and for all spatially smoothed image as the input for computing filter bank responses for all the possible orientations. In the case, when the underlying scale-space representation is instead based on convolutions with affine Gaussian kernels, we do, of course, have to redo the underlying spatial smoothing operation for each combination of the parameters $\sigma_{1}$, $\sigma_{2}$ and $\varphi$. We can, however, nevertheless reuse the same underlying spatially smoothed image for all the combinations of the orders $m_{1}$ and $m_{2}$ of spatial differentiation.
@@ 5.2 Method for defining discrete directional derivative approximation masks
To define a discrete derivative approximation mask $\delta_{\varphi^{m_{1}}\bot\varphi^{m_{2}}}$, for computing an approximation of the directional derivative $L_{\varphi^{m_{1}}\bot\varphi^{m_{2}}}$ from an already smoothed scale-space representation $L$ according to
@@ 6 Summary and conclusions
Regarding the purely spatial smoothing operation, the discrete analogue of the Gaussian kernel stands out as having the best theoretical properties over the entire scale range, from scale levels approaching zero to coarser scales. The results obtained from the sampled Gaussian kernel may deviate substantially from their continuous counterparts, when the scale parameter $\sigma$ is less than about 0.5 or 0.75. For $\sigma$ greater than about 1, the sampled Gaussian kernel does, on the other hand, lead to numerically very good approximations of results obtained from the corresponding continuous theory.
@@
The fact that the sampled Gaussian derivative kernels for sufficiently coarse scales lead to the closest approximations of the corresponding fully continuous theory should, however, not preclude from basing the analysis on the discrete analogues of Gaussian derivatives at coarser scales. If necessary, deviations between the results obtained from the discrete analogues of Gaussian derivatives and the corresponding fully continuous theory can, in principle, be compensated for by complementary calibration procedures, or by deriving corresponding genuinely discrete analogues of the relevant entities in the analysis. Additionally, in situations when a larger number of Gaussian derivative responses are to be computed simultaneously, this can be accomplished with substantially higher computational efficiency, if basing the scale-space analysis on the discrete analogue of the Gaussian kernel, which only involves a single spatial smoothing stage of large spatial support, from which each derivative approximation can then be computed using a small-support central difference operator.
As a complement to the presented methodologies of discretizing Gaussian smoothing and Gaussian derivative computations, we have also in Section 5 presented a computationally very efficient ways of computing directional derivatives of different orders and of different orientations, which is highly useful for computing filter bank type responses for different purposes in computer vision. When using the discrete analogue of the Gaussian kernel for smoothing, the presented discrete directional derivative approximation masks can be applied at any scale. If using either sampled Gaussian kernels or integrated Gaussian kernels for spatial smoothing, including extensions of rotationally symmetric kernels to anisotropic affine Gaussian kernels, the discrete derivative approximation masks can be used, provided that the scale parameter is sufficiently large in relation to the desired accuracy of the resulting numerical approximation.